OpenStudy (anonymous):

What is the answer 2/4 × 2

Vocaloid (vocaloid):

2/4 = 1/2 1/2 * 2 = 1

OpenStudy (anonymous):

Thanks!

OpenStudy (photon336):

@RoseyPosey technically this is the multiplicative inverse if you multiply any number by 1/itself you get one see below. \[x*(\frac{ 1 }{ x }) = \frac{ x }{ x } = 1\]

OpenStudy (photon336):

\[2*\frac{ 1 }{ 2 } = \frac{ 2 }{ 2 } = 1\]

OpenStudy (anonymous):

1/2*2=1 ,,,,,2 and 1/2 are reciprocals...